OpenStudy (anonymous):

x^4-144=0 solved using the quadratic formula

mathslover (mathslover):

I will write it as : \(x^2 + 0(x) - 144 = 0\) Can you predict what are a, b and c here? a = coefficient of x^2 b = coefficient of x c = constant term

OpenStudy (anonymous):

put x^2=t \[t^2+0t-144=0,solve ~for ~t\] then again replace t by x^2.